OpenStudy (anonymous):

G is the incenter of Δ ABC. m∠A = 55 m∠B = 70 m∠C = 55 What is the m∠ABF? 25 35 27.5 40

OpenStudy (anonymous):

where is point F? or did you mean ABG?

OpenStudy (anonymous):

 Angle ABF = half of angle ABC = 1/2 x 70 = 35 Angle BAF = half of angle BAC = 1/2 x 55 = 27.5 Sum of angles in triangle ABF = 180 = 35 + 27.5 + ABF ABF = 180 - 35 - 27.5 = ?

OpenStudy (anonymous):

An incenter is made up of three Angle Bisectors so each Angle is cut in half so ABG is refrerring to angle B because B is in the middle of the three letters representing points. so your answer is half of angle B which you said is 70 so your answer should be 35 degrees

OpenStudy (anonymous):

Oh so point F is not G

OpenStudy (anonymous):

But I believe my answer is still correct the degree of angle ABF is 35

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

I just learned this Im happy to help!